OpenStudy (dr.gray):

which equation represents the cross product equation derived from the proportion 6 over x =15 over 10 A. 6 x =150 B. 10 x=90 C. 15 x = 60 D.15 x = 610

OpenStudy (dr.gray):

the only reason why i asked this question is because i suck at math

OpenStudy (serenity74):

6/x = 15/10 sp cross multiply would be x * 15 = 6* 10 so 15 x = 60 is the answer

OpenStudy (dr.gray):

thx

OpenStudy (serenity74):

:))

OpenStudy (serenity74):

If you need anymore help let me know! :D

OpenStudy (dr.gray):

ok i'm about to post another question

OpenStudy (serenity74):

Alright just tag me and I will try my best to answer it :)

OpenStudy (dr.gray):

how do i tag you

OpenStudy (serenity74):

just do @Serenity74 and bam you got me

OpenStudy (mathmale):

Please delete this entry: "the only reason why i asked this question is because i suck at math." In its place, write, "I struggle with math sometimes, but am sure that with enough study and practice I'll be fine with it."

OpenStudy (dangalique):

I think that its D) 15x=610

OpenStudy (serenity74):

@dangalique how so

OpenStudy (dangalique):

It seems to be the only one that can work!!! : )

OpenStudy (mathmale):

Yes, Dr.Gray, it's important that you show your work, so that we can give you feedback on it.

OpenStudy (mathmale):

Dr.G: what does "cross-product" mean?

OpenStudy (serenity74):

but 6 * 10 doesn't = 610....

OpenStudy (dangalique):

i sent u an email mathmale. : )